You purchase a 7% annual coupon bond for $871.65 today. You then sell the bond after
holding it for one year and earn a 9% return on your investment. What is the price for
which you would have sold the bond?
Answer #1

F = P *(1+i)^t

i = 9% = 0.09

t = 1 yr

P = 871.65

Sale value of bond = 871.65 * (1+0.09) = 871.65 * 1.09 = 950.0985 ~ 950.10